# Kullback–Leibler Potential for Non-Ergodic Replication Dynamics: An Information-Theoretic Second Law

### v1

@@ III.2 DNA replication as a block-diagonal substitution process
$h_{b}(\varepsilon)$ $=-\varepsilon\log_{2}\varepsilon-(1-\varepsilon)\log_{2}(1-\varepsilon).$
Readouts and figures.
The trajectories in Fig.2 show an increase in $H$ and $H_{\times}$, a decrease in $D_{\mathrm{KL}}(p_{n}\|q_{n})$, and a monotone decay of $V(p_{n})$, in agreement with Theorem 1. Figure 3 shows the potential landscape $V(p)=w_{1}V_{1}(x)+w_{2}V_{2}(y)$ on $(x,y)$, where $x$ is the A fraction in $\{\mathrm{A},\mathrm{T}\}$ and $y$ the C fraction in $\{\mathrm{C},\mathrm{G}\}$.
Figure 2: DNA block-diagonal replication: time series of informational metrics (nats).Shannon entropy, Cross entropy, KLD, and KLD potential. We iterate a four-state Markov chain on $\{\mathrm{A},\mathrm{T},\mathrm{C},\mathrm{G}\}$ with a block-diagonal kernel $\widetilde{T}=\mathrm{diag}(T_{1},T_{2})$ for $N=50$ steps. Base rates: $\alpha=0.020$, $\beta=0.010$, $\gamma=0.015$, $\delta=0.015$. Proofreading channel $\mathcal{R}$ with $\alpha^{\prime}=0.005$, $\beta^{\prime}=0.003$, $\gamma^{\prime}=\delta^{\prime}=0.004$ is mixed at $\rho=0.30$, giving effective rates $\alpha_{\!e}=0.0155$, $\beta_{\!e}=0.0079$, $\gamma_{\!e}=\delta_{\!e}=0.0117$ and block invariants $\pi_{1}^{\ast}=(0.3376,\,0.6624)$, $\pi_{2}^{\ast}=(0.5,\,0.5)$. Initial distribution $p_{0}=(0.6,0.1,0.2,0.1)$ so that $w_{1}(p_{0})=0.7$, $w_{2}(p_{0})=0.3$. At each step $q_{n}=p_{n}\widetilde{T}$ and we record: top-left, Shannon entropy $H(q_{n})$; top-right, cross-entropy $H(p_{n},q_{n})$; bottom-left, step divergence $D_{\mathrm{KL}}(p_{n}\|q_{n})$; bottom-right, KLD potential $V(p_{n})=w_{1}(p_{0})\,D_{\mathrm{KL}}(p_{n}^{(1)}\|\pi_{1}^{\ast})+w_{2}(p_{0})\,D_{\mathrm{KL}}(p_{n}^{(2)}\|\pi_{2}^{\ast})$. As replication proceeds, $H$ and $H_{\times}$ increase, while both $D_{\mathrm{KL}}(p_{n}\|q_{n})$ and $V(p_{n})$ decrease monotonically; the monotonicity of $V$ follows from Theorem 1.
The contour plot shows the KL potential $V(p)=w_{1}D_{\mathrm{KL}}\!\bigl{(}p^{(1)}\!\parallel\!\pi_{1}^{\ast}\bigr{)}+w_{2}D_{\mathrm{KL}}\!\bigl{(}p^{(2)}\!\parallel\!\pi_{2}^{\ast}\bigr{)}$ over $(x,y)\in[0,1]^{2}$, where $x$ is the A fraction in the AT block and $y$ is the C fraction in the GC block. The unique minimum (blue marker) occurs at $(x^{\ast},y^{\ast})=(\pi_{1}^{\ast}(\mathrm{A}),\pi_{2}^{\ast}(\mathrm{C}))\approx(0.338,0.600)$, i.e., at the block-wise invariants determined by the effective rates. Because $V$ is a sum of blockwise KL terms, the level sets are nearly axis-aligned ellipses (here close to circular since $w_{1}=w_{2}=0.5$ and the curvatures are similar). A second-order expansion around $(x^{\ast},y^{\ast})$ gives $V(x,y)\approx\tfrac{w_{1}}{2\,\mu(1-\mu)}(x-\mu)^{2}+\tfrac{w_{2}}{2\,\nu(1-\nu)}(y-\nu)^{2}$ with $\mu=\pi_{1}^{\ast}(\mathrm{A})\approx 0.338$ and $\nu=\pi_{2}^{\ast}(\mathrm{C})=0.600$, yielding coefficients $\simeq 1.12$ and $\simeq 1.04$ (nats), respectively—consistent with the nearly isotropic contours. Increasing proofreading bias decreases $\mu(1-\mu)$ or $\nu(1-\nu)$, steepening the basin and accelerating the monotonic decay of $V(p_{n})$ (Theorem 1). The separable, convex landscape highlights that, under block invariance, the non-ergodic relaxation proceeds independently within blocks and converges to the reachable steady point $(x^{\ast},y^{\ast})$(Fig. 3).
Figure 3: KLD potential landscape $V(p)=w_{1}D_{\mathrm{KL}}(p^{(1)}\parallel\pi_{1}^{\ast})+w_{2}D_{\mathrm{KL}}(p^{(2)}\parallel\pi_{2}^{\ast})$ (nats) over $(x,y)\in[0,1]^{2}$, where $x$ is the A fraction in the AT block and $y$ is the C fraction in the GC block. Proofreading mixture $\rho=0.30$; AT rates $\alpha=0.020$, $\beta=0.010$, $\alpha^{\prime}=0.005$, $\beta^{\prime}=0.003$ give $\alpha_{\mathrm{eff}}=0.0155$, $\beta_{\mathrm{eff}}=0.0079$ and $\pi_{1}^{\ast}=(\beta_{\mathrm{eff}}/(\alpha_{\mathrm{eff}}+\beta_{\mathrm{eff}}),,\alpha_{\mathrm{eff}}/(\alpha_{\mathrm{eff}}+\beta_{\mathrm{eff}}))\approx(0.338,,0.662)$. GC rates (asymmetric) $\gamma=0.014$, $\delta=0.021$, $\gamma^{\prime}=0.004$, $\delta^{\prime}=0.006$ give $\gamma_{\mathrm{eff}}=0.0110$, $\delta_{\mathrm{eff}}=0.0165$ and $\pi_{2}^{\ast}=(\delta_{\mathrm{eff}}/(\gamma_{\mathrm{eff}}+\delta_{\mathrm{eff}}),,\gamma_{\mathrm{eff}}/(\gamma_{\mathrm{eff}}+\delta_{\mathrm{eff}}))\approx(0.60,,0.40)$. Block weights $w_{1}=w_{2}=0.5$. The blue marker indicates $(x^{\ast},y^{\ast})=(\pi_{1}^{\ast}(\mathrm{A}),,\pi_{2}^{\ast}(\mathrm{C}))\approx(0.34,,0.60)$. Grid $101\times 101$, colormap viridis.
An information-theoretic second law (minimal form)
State function and step variables.
Let
$\Delta V_{n}:=V(p_{n})-V(p_{n+1}),$ (41)
and
$D_{n}:=D_{\mathrm{KL}}(p_{n}\|q_{n}),\qquad q_{n}:=p_{n}T.$ (42)
By Theorem 1 and nonnegativity of KLD, one has $\Delta V_{n}\geq 0$ and $D_{n}\geq 0$.
Step production (dimensionless) and its sum.
Define the per-step production
$\mathcal{S}_{n}:=D_{n}+\Delta V_{n},$ (43)
and the cumulative production over $n=0,\dots,N-1$,
$\mathcal{S}^{[0:N)}:=\sum_{n=0}^{N-1}\mathcal{S}_{n}=\sum_{n=0}^{N-1}D_{n}\;+\;V(p_{0})-V(p_{N}).$ (44)
In physical units,
$\sigma_{\rm rep}^{(n)}:=k_{\mathrm{B}}T\,\mathcal{S}_{n},\qquad\Sigma_{\rm rep}^{(N)}:=k_{\mathrm{B}}T\,\mathcal{S}^{[0:N)}.$ (45)
Theorem 2 (Second law for non-ergodic replication: minimal form).
For any block-invariant trajectory $\{p_{n}\}_{n=0}^{N}$,
$\mathcal{S}_{n}\geq 0\quad(n=0,1,\dots,N-1),$ (46)
hence
$\mathcal{S}^{[0:N)}\geq 0.$ (47)
Equivalently, $\sigma_{\rm rep}^{(n)}\geq 0$ and $\Sigma_{\rm rep}^{(N)}\geq 0$.
Idea.
Immediate from $D_{n}\geq 0$ and $\Delta V_{n}\geq 0$. ∎
Remark 2 (Equality and strictness).
$\mathcal{S}_{n}=0$ iff $D_{n}=\Delta V_{n}=0$; otherwise $\mathcal{S}_{n}>0$.
IV Discussion
@@
The minimal second law above decomposes the step production into two nonnegative parts, $D_{n}$ and $\Delta V_{n}$, cf. (43). In physical units, the potential drop defines an informational free-energy change
@@
If the reachable steady set $\Pi(p_{0})$ collapses to a single invariant distribution (ergodic limit), $V$ reduces to the standard KL-to-steady Lyapunov function. Under non-ergodic, block-invariant constraints, $V$ is the minimal KL distance to the steady set $\Pi(p_{0})$, so its decay quantifies relaxation under conserved coarse masses. This distinction is essential for replication processes where structural constraints inhibit global mixing.
@@
Equality in (17) requires blockwise tightness of the data-processing inequality (Remark 1); operationally, this corresponds to being at (or effectively on) a block-wise steady state. Away from such fixed points, the decrease is strict and $V$ serves as a sensitive progress variable for replication-driven relaxation.
V Conclusion and Outlook
@@
Appendix A Simulation protocol for Fig. 1 (Gaussian-copy model)
@@
Domain and initialization. We use a $256\times 256$ pixel grid. The image domain is partitioned into a regular $B_{x}\times B_{y}$ tiling of rectangular blocks (Fig. 1 uses $B_{x}=B_{y}=4$; we also verified $2\times 2$ and $128\times 128$). Inside each block, a binary pattern is drawn i.i.d. from a Bernoulli law with a checkerboard success probability: even (i+j)(i{+}j) blocks use $0.8$, odd (i+j)(i{+}j) blocks use $0.2$. The random seed is fixed to $7$. The resulting array is normalized to a probability mass function (pmf) $p_{0}=I_{0}/\!\sum_{x,y}I_{0}(x,y)$.
@@
  • Ergodic (global): Gaussian smoothing is applied to the whole image to produce $q_{n}$.
@@
In both cases we update $p_{n+1}=q_{n}$. We run $n_{\text{steps}}=50$ and store snapshots at $n\in\{0,10,20,30,40,50\}$.
@@
KLD potential $V(p_{n})$. We use the Lyapunov-type potential defined consistently with the main text:
  • Ergodic case: $V(p_{n})=D_{\mathrm{KL}}(p_{n}\|u)$, where $u$ is the uniform pmf over all $256\times 256$ pixels.
  • Non-ergodic (blockwise) case: Let $b$ index blocks, $w_{b}=\sum_{(x,y)\in b}p_{n}(x,y)$ the block mass, $p_{b}$ the within-block normalized pmf, and $u_{b}$ the blockwise uniform pmf. Then
    $V(p_{n})=\sum_{b}w_{b}\,D_{\mathrm{KL}}\!\bigl{(}p_{b}\|u_{b}\bigr{)}.$ (52)
This definition matches the theoretical reachable-steady-set construction and yields the monotone decay predicted by Theorem 1.
Visualization and units. Time-series plots show $H$, $H(p_{n},q_{n})$, $D_{\mathrm{KL}}(p_{n}\|q_{n})$, and $V(p_{n})$ versus the replication step $n$. Snapshots at the specified steps use a common grayscale colorbar. All values are reported in bits.
Sanity condition on $\sigma$. If the Gaussian width exceeds roughly one quarter of the smallest block edge, boundary effects may dominate. In Fig. 1 (block size $64{\times}64$, $\sigma=1.5$) this issue does not arise. The same qualitative trends were observed for $2{\times}2$ and $128{\times}128$ partitions (not shown).
Appendix B: Simulation details for Fig. 2 (time series)
B.1 Model and parameters.
We consider a block-diagonal, discrete-time Markov kernel on $\mathcal{X}=\{\mathrm{A},\mathrm{T},\mathrm{C},\mathrm{G}\}$ with blocks $\mathcal{X}_{1}=\{\mathrm{A},\mathrm{T}\}$ and $\mathcal{X}_{2}=\{\mathrm{C},\mathrm{G}\}$. One replication step uses an extension channel $T$ mixed with a block-wise proofreading/repair channel $\mathcal{R}$ with probability $\rho$:
$\widetilde{T}\;=\;(1-\rho)\,T\;+\;\rho\,\mathcal{R}\;=\;\mathrm{diag}\!\bigl{(}T_{1}^{\mathrm{eff}},T_{2}^{\mathrm{eff}}\bigr{)}.$
Each $2\times 2$ block takes the form
$T_{1}=\begin{pmatrix}1-\alpha&\alpha\\[2.0pt] \beta&1-\beta\end{pmatrix},\quad T_{2}=\begin{pmatrix}1-\gamma&\gamma\\[2.0pt] \delta&1-\delta\end{pmatrix}$,
$R_{1}=\begin{pmatrix}1-\alpha^{\prime}&\alpha^{\prime}\\[2.0pt] \beta^{\prime}&1-\beta^{\prime}\end{pmatrix},\quad R_{2}=\begin{pmatrix}1-\gamma^{\prime}&\gamma^{\prime}\\[2.0pt] \delta^{\prime}&1-\delta^{\prime}\end{pmatrix}$.
Effective (mixed) rates are
$\alpha_{\!\mathrm{eff}}=(1-\rho)\alpha+\rho\alpha^{\prime},\ \beta_{\!\mathrm{eff}}=(1-\rho)\beta+\rho\beta^{\prime},$
$\gamma_{\!\mathrm{eff}}=(1-\rho)\gamma+\rho\gamma^{\prime},\ \delta_{\!\mathrm{eff}}=(1-\rho)\delta+\rho\delta^{\prime}.$
Thus
$T_{1}^{\mathrm{eff}}=\begin{pmatrix}1-\alpha_{\!\mathrm{eff}}&\alpha_{\!\mathrm{eff}}\\[2.0pt] \beta_{\!\mathrm{eff}}&1-\beta_{\!\mathrm{eff}}\end{pmatrix},\quad T_{2}^{\mathrm{eff}}=\begin{pmatrix}1-\gamma_{\!\mathrm{eff}}&\gamma_{\!\mathrm{eff}}\\[2.0pt] \delta_{\!\mathrm{eff}}&1-\delta_{\!\mathrm{eff}}\end{pmatrix}$.
Block-wise invariants are
$\pi_{1}^{\ast}=\!\Bigl{(}\tfrac{\beta_{\!\mathrm{eff}}}{\alpha_{\!\mathrm{eff}}+\beta_{\!\mathrm{eff}}},\tfrac{\alpha_{\!\mathrm{eff}}}{\alpha_{\!\mathrm{eff}}+\beta_{\!\mathrm{eff}}}\Bigr{)}\approx(0.3376,\,0.6624),$
$\pi_{2}^{\ast}=\!\Bigl{(}\tfrac{\delta_{\!\mathrm{eff}}}{\gamma_{\!\mathrm{eff}}+\delta_{\!\mathrm{eff}}},\tfrac{\gamma_{\!\mathrm{eff}}}{\gamma_{\!\mathrm{eff}}+\delta_{\!\mathrm{eff}}}\Bigr{)}=(0.5,\,0.5).$
B.2 Initial condition and conserved masses.
@@
B.3 Recorded quantities (natural logs; nats).
@@
The KLD potential (Def. (11)) reduces to the block sum
@@
B.5 Minimal Python excerpt (reproducibility).
import numpy as np
def H_shannon(p):
    m = p > 0
    return -np.sum(p[m]*np.log(p[m]))
def H_cross(p, q):
    m = (p > 0) & (q > 0)
    return -np.sum(p[m]*np.log(q[m]))
def Dkl(p, q):
    m = (p > 0) & (q > 0)
    return np.sum(p[m]*(np.log(p[m]) - np.log(q[m])))
def V_block2(p, w1_0, w2_0, pi1, pi2):
    pA,pT,pC,pG = p
    w1, w2 = pA+pT, pC+pG
    p1 = np.array([pA, pT])/w1
    p2 = np.array([pC, pG])/w2
    return w1_0*Dkl(p1, pi1) + w2_0*Dkl(p2, pi2)
# parameters (nats)
alpha, beta   = 0.020, 0.010
gamma, delta  = 0.015, 0.015
alpha_p, beta_p = 0.005, 0.003
gamma_p, delta_p = 0.004, 0.004
rho = 0.30
ae = (1-rho)*alpha + rho*alpha_p
be = (1-rho)*beta  + rho*beta_p
ge = (1-rho)*gamma + rho*gamma_p
de = (1-rho)*delta + rho*delta_p
T1 = np.array([[1-ae, ae],[be, 1-be]], float)
T2 = np.array([[1-ge, ge],[de, 1-de]], float)
T  = np.zeros((4,4), float); T[:2,:2]=T1; T[2:,2:]=T2
pi1 = np.array([be/(ae+be), ae/(ae+be)], float)
pi2 = np.array([de/(ge+de), ge/(ge+de)], float)
p = np.array([0.6,0.1,0.2,0.1], float); p = p/np.sum(p)
w1_0, w2_0 = p[0]+p[1], p[2]+p[3]
N = 50
Hq, Hx, D, V = [], [], [], []
for n in range(N+1):
    V.append(V_block2(p, w1_0, w2_0, pi1, pi2))   # V(p_n)
    if n == N: break
    q = p @ T # q_n = p_n T
    Hq.append(H_shannon(q))                       # H(q_n)
    Hx.append(H_cross(p, q))                      # H(p_n, q_n)
    D.append(Dkl(p, q))                           # D_KL(p_n || q_n)
    p = q
Appendix C: Simulation details for Fig. 3 (potential landscape)
C.1 Model and objective.
We consider two invariant blocks $\mathcal{X}_{1}=\{\mathrm{A},\mathrm{T}\}$ and $\mathcal{X}_{2}=\{\mathrm{C},\mathrm{G}\}$. On the grid $(x,y)\in[0,1]^{2}$, we parameterize a block-wise distribution
@@
where $x$ is the A-fraction in the AT block and $y$ is the C-fraction in the GC block. The KL potential (Def. (11)) reduces to the separable sum
@@
C.2 Effective rates and steady states.
@@
AT block (Fig. 2): $\alpha=0.020,\ \beta=0.010,\ \alpha^{\prime}=0.005,\ \beta^{\prime}=0.003$ $\Rightarrow\ \alpha_{\!e}=0.0155,\ \beta_{\!e}=0.0079$. GC block: $\gamma=0.014,\ \delta=0.021,\ \gamma^{\prime}=0.004,\ \delta^{\prime}=0.006$ $\Rightarrow\ \gamma_{\!e}=0.0110,\ \delta_{\!e}=0.0165$. The block-wise invariants (for [[1a,a],[b,1b]][[1-a,a],[b,1-b]]) are
@@
C.4 Minimal Python excerpt (reproducibility).
@@
# GC block (asymmetric, to place y*  0.60)

### v2

@@ 2.2 DNA replication as a block-diagonal substitution process
Summary.
We (i) defined $D_{n}$ and $\Delta V_{n}$, (ii) separated their sources of nonnegativity (KLD vs. Lyapunov monotonicity), (iii) derived the telescoping-form inequality, and (iv) characterized equality, interpretation, and extensions. This mirrors Hatano–Sasa/Esposito–Van den Broeck–type formulations for Markov jump processes while crucially referencing a set of reachable steady states rather than a single invariant measure.
3 Discussion
@@
The minimal second law above decomposes the step production into two nonnegative parts, $D_{n}$ and $\Delta V_{n}$, cf. (45). In physical units, the potential drop defines a informational free-energy change.
@@
If the reachable steady set $\Pi(p_{0})$ collapses to a single invariant distribution (ergodic limit), $V$ reduces to the standard KL-to-steady Lyapunov function. Under non-ergodic block-invariant constraints, $V$ is minimal KLD to the steady set $\Pi(p_{0})$, so its decay quantifies relaxation under conserved coarse masses. This distinction is essential for replication processes where structural constraints inhibit global mixing.
@@
The equality in (14) requires blockwise tightness of the data processing inequality (Remark 1); operationally, this corresponds to being in (or effectively in) a blockwise steady state. Despite such fixed points, the decrease is strict and $V$ serves as a sensitive progress variable for replication-driven relaxation.
Thermodynamic meaning of a non-zero asymptotic $V$.
Interpretation at long times. The following clarifies the thermodynamic reading of a non-zero asymptotic $V$. Our KLD potential $V$ quantifies distance to the reachable steady set that preserves coarse variables (block masses). In a driven NESS with primitive blocks, $V(p_{n})\!\to\!0$ even though steady dissipation (housekeeping heat) can remain positive; $V$ is a state function, not a direct measure of instantaneous entropy production. A strictly positive long-time value emerges only when the trajectory is persistently kept away from the reachable set, e.g. due to weak inter-block leakage, non-primitive blocks, or active enforcement of coarse masses. In that case, choosing $\pi_{j}=p^{(j)}$ yields
$V(p)=\sum_{j}w_{j}(p)\,\log\!\frac{w_{j}(p)}{w_{j}(p_{0})},$
so $k_{\mathrm{B}}T\,V(p)$ is the minimal informational free-energy cost to refresh the state back to the admissible coarse constraint. If refreshes occur with frequency $f$, the corresponding maintenance power satisfies $\mathcal{P}_{\min}\geq f\,k_{\mathrm{B}}T\,V_{\!*}$, where $V_{\!*}$ is the stationary value under the leak/drive. Thus, a non-zero asymptotic $V$ can be read as a lower bound on the energetic cost of maintaining a coarse-grained steady state while microscopic erasure/copying continuously occurs [15, 16].
4 Conclusion and Outlook
@@
Appendix A: Derivation and leakage threshold
Closed-form expression for $V_{\delta}$ (corrected).
Write $\pi=\sum_{j=1}^{m}w_{j}\,\pi_{j}$ with $w\in[0,1]^{m}$, $\sum_{j=1}^{m}w_{j}=1$ and $\pi_{j}\in\Delta(\mathcal{X}_{j})$. By the block decomposition of KL,
$D_{\mathrm{KL}}(p\|\pi)=\sum_{j=1}^{m}w_{j}(p)\,D_{\mathrm{KL}}\!\bigl(p^{(j)}\|\pi_{j}\bigr)+\sum_{j=1}^{m}w_{j}(p)\,\log\!\frac{w_{j}(p)}{w_{j}}.$
For fixed $w$, the inner minimization is attained at $\pi_{j}=p^{(j)}$. Thus
$V_{\delta}(p)=\min_{\begin{subarray}{c}w:\ \sum_{j=1}^{m}w_{j}=1\\ a_{j}\leq w_{j}\leq b_{j}\end{subarray}}\;\varphi(w),\qquad\varphi(w):=\sum_{j=1}^{m}w_{j}(p)\,\log\!\frac{w_{j}(p)}{w_{j}},$
with $a_{j}:=w_{j}(p_{0})-\delta$, $b_{j}:=w_{j}(p_{0})+\delta$. This is a convex program. The Lagrangian with multipliers $\tau\in\mathbb{R}$ (simplex) and $\mu_{j}^{\pm}\geq 0$ (box) is
$\mathcal{L}(w,\tau,\mu^{\pm})=\sum_{j=1}^{m}w_{j}(p)\,\log\frac{w_{j}(p)}{w_{j}}+\tau\!\left(\sum_{j=1}^{m}w_{j}-1\right)+\sum_{j=1}^{m}\mu_{j}^{+}(w_{j}-b_{j})+\sum_{j=1}^{m}\mu_{j}^{-}(a_{j}-w_{j}).$
KKT conditions (necessary and sufficient here) yield, for an optimum $(w^{\star},\tau^{\star},\mu^{\pm\star})$,
$\frac{w_{j}(p)}{w_{j}^{\star}}+\tau^{\star}+\mu_{j}^{+\star}-\mu_{j}^{-\star}=0,\quad\sum_{j=1}^{m}w_{j}^{\star}=1,\quad a_{j}\leq w_{j}^{\star}\leq b_{j},$
$\mu_{j}^{+\star}(w_{j}^{\star}-b_{j})=\mu_{j}^{-\star}(a_{j}-w_{j}^{\star})=0,\qquad\mu_{j}^{\pm\star}\geq 0.$
Hence interior coordinates ($a_{j}<w_{j}^{\star}<b_{j}$) satisfy $w_{j}^{\star}=w_{j}(p)/\tau^{\star}$; active coordinates stick to the nearest bound. Equivalently, the solution is the scaled-and-clipped vector that also satisfies the simplex constraint:
$w_{j}^{\star}=\Bigl[\frac{w_{j}(p)}{\tau^{\star}}\Bigr]_{[\,a_{j},\,b_{j}\,]},\qquad\sum_{j=1}^{m}w_{j}^{\star}=1,$
where $\tau^{\star}>0$ is uniquely determined by the normalization. Substituting back gives
$\boxed{\;V_{\delta}(p)=\sum_{j=1}^{m}w_{j}(p)\,\log\!\frac{w_{j}(p)}{\,w_{j}^{\star}\,},\quad w_{j}^{\star}=\Bigl[\frac{w_{j}(p)}{\tau^{\star}}\Bigr]_{[\,w_{j}(p_{0})-\delta,\;w_{j}(p_{0})+\delta\,]},\ \ \sum_{j}w_{j}^{\star}=1.\;}$ (39)
The associated minimizer is $\pi_{j}^{\star}=p^{(j)}$ on each block, i.e. $\pi^{\star}\!\restriction_{\mathcal{X}_{j}}=w_{j}^{\star}\,p^{(j)}$.
Continuity.
As $\delta\to 0$, the box $a_{j},b_{j}]$ collapses to $w_{j}(p_{0})$, so $w^{\star}\to w(p_{0})$ and $V_{\delta}(p)\to V(p)$.
Sufficient threshold for strict decrease under leakage.
Let $w_{n}=(w_{1}(p_{n}),\dots,w_{m}(p_{n}))$ and $\Delta w:=w_{n+1}-w_{n}$. Define
$\kappa_{n}:=\sum_{j=1}^{m}w_{j}(p_{0})\Bigl[D_{\mathrm{KL}}\!\bigl(p_{n}^{(j)}\|\pi_{j}^{\ast}\bigr)-D_{\mathrm{KL}}\!\bigl(T_{j}p_{n}^{(j)}\|\pi_{j}^{\ast}\bigr)\Bigr]\ \geq 0,\qquad w_{\min}:=\min_{1\leq j\leq m}w_{j}(p_{0}).$
Using the KL block decomposition between steps,
$V(p_{n+1})-V(p_{n})$ $=\sum_{j=1}^{m}\!\Bigl[w_{j}(p_{n+1})-w_{j}(p_{n})\Bigr]\!\log\!\frac{w_{j}(p_{n+1})}{w_{j}(p_{0})}$
$\quad+\sum_{j=1}^{m}w_{j}(p_{0})\!\Bigl[D_{\mathrm{KL}}\!\bigl(T_{j}p_{n}^{(j)}\|\pi_{j}^{\ast}\bigr)-D_{\mathrm{KL}}\!\bigl(p_{n}^{(j)}\|\pi_{j}^{\ast}\bigr)\Bigr],$
the second line is $\kappa_{n}$. By convexity of $x\mapsto x\log x$ and $w_{j}\geq w_{\min}$, the first line is bounded by $\tfrac{1}{2w_{\min}}\lVert\Delta w\rVert_{1}^{2}$. Therefore
$V(p_{n+1})-V(p_{n})\ \leq\ \frac{1}{2w_{\min}}\lVert\Delta w\rVert_{1}^{2}-\kappa_{n},$
and a sufficient condition for a strict one-step decrease is
$\boxed{\;\lVert\Delta w\rVert_{1}<2\sqrt{\,w_{\min}\,\kappa_{n}\,}\ \Rightarrow\ V(p_{n+1})<V(p_{n})\!.\;}$
Appendix B:Simulation protocol for Fig. 1 (Gaussian-copy model)
@@
Domain and initialization. We use three types of pixel grid. The image domain is partitioned into a regular $B_{x}\times B_{y}$ tiling of rectangular blocks (Fig. 1 uses $B_{x}=B_{y}=4$; we also verified $2\times 2$, $4\times 4$ and $128\times 128$). Inside each block, a binary pattern is drawn i.i.d. from a Bernoulli law with a checkerboard success probability: even (i+j)(i{+}j) blocks use $0.8$, odd (i+j)(i{+}j) blocks use $0.2$. The random seed is fixed to $7$. The resulting matrix is normalized to a probability mass function (pmf) $p_{0}=I_{0}/\!\sum_{x,y}I_{0}(x,y)$.
@@
  • Ergodic (global): Gaussian smoothing is applied to the entire image to produce $q_{n}$.
@@
In both cases, we update $p_{n+1}=q_{n}$. We run $n_{\text{steps}}=50$ and store snapshots at $n\in\{0,10,20,30,40,50\}$.
@@
KLD potential $V(p_{n})$. We use the Lyapunov-type potential consistent with the main text:
  • Ergodic (global) case: $V(p_{n})=D_{\mathrm{KL}}(p_{n}\|\pi^{\ast})$, where $\pi^{\ast}$ is the unique invariant histogram under global convolution dynamics (approximated numerically by iterating $T_{\sigma}$ to stationarity).
  • Non-ergodic (blockwise) case: Let $b$ index blocks, $w_{b}=\sum_{(x,y)\in b}p_{n}(x,y)$, $p_{b}$ the normalized histogram within the block, and $u_{b}$ the uniform histogram over block $b$. Then
    $V(p_{n})=\sum_{b}w_{b}\,D_{\mathrm{KL}}\!\bigl(p_{b}\|u_{b}\bigr),$
    which coincides with the distance to the reachable steady set with fixed block masses.
Visualization and units. Time series plots show $H$, $H(p_{n},q_{n})$, $D_{\mathrm{KL}}(p_{n}\|q_{n})$, and $V(p_{n})$ versus the replication step $n$. Snapshots at the specified steps use a common grayscale colorbar. All values are reported in bits.
Sanity condition on $\sigma$. If the Gaussian width exceeds approximately one-quarter of the smallest block edge, boundary effects may dominate. In Fig. 1 (block size $128{\times}128$, $\sigma=1.5$) this issue does not arise. The same qualitative trends were observed for $2{\times}2$ and $4{\times}4$ partitions.
Appendix C:Simulation details for Fig. 2 (time series)
Model reference & parameters (Fig. 2).
We use the two-block Markov model of App. C.2 (block-diagonal kernel; one-step mixture $\widetilde{T}=(1-\rho)T+\rho\mathcal{R}$; effective rates as in Eq. (43) and blockwise invariants $\pi_{j}^{\ast}$ therein). For Fig. 2 we set (nats)
$\alpha=0.020,\ \beta=0.010,\ \gamma=\delta=0.015,\quad\alpha^{\prime}=0.005,\ \beta^{\prime}=0.003,\ \gamma^{\prime}=\delta^{\prime}=0.004,\quad\rho=0.30,$
giving
$\alpha_{\!e}=0.0155,\ \beta_{\!e}=0.0079,\ \gamma_{\!e}=\delta_{\!e}=0.0117,$
and invariants
$\pi_{1}^{\ast}=\Bigl(\tfrac{\beta_{\!e}}{\alpha_{\!e}+\beta_{\!e}},\,\tfrac{\alpha_{\!e}}{\alpha_{\!e}+\beta_{\!e}}\Bigr)\approx(0.3376,\,0.6624),\qquad\pi_{2}^{\ast}=(0.5,\,0.5).$
Initial condition and conserved masses.
@@
Recorded quantities (natural logs; nats).
@@
The KLD potential (Def. (1)) reduces to the sum of blocks.
@@
Appendix D: Simulation details for Fig. 3 (potential landscape)
D.1 Model and objective.
We consider two invariant blocks $\mathcal{X}_{1}=\{\mathrm{A},\mathrm{T}\}$ and $\mathcal{X}_{2}=\{\mathrm{C},\mathrm{G}\}$. On the grid $(x,y)\in[0,1]^{2}$, we parameterize a block-wise distribution.
@@
where $x$ is the A fraction in the AT block and $y$ is the C fraction in the GC block. The KLD potential (Def. (1)) reduces to the separable sum
@@
D.2 Effective rates and steady states.
@@
AT block (Fig. 2): $\alpha=0.020,\ \beta=0.010,\ \alpha^{\prime}=0.005,\ \beta^{\prime}=0.003$ $\Rightarrow\ \alpha_{\!e}=0.0155,\ \beta_{\!e}=0.0079$. GC block: $\gamma=0.014,\ \delta=0.021,\ \gamma^{\prime}=0.004,\ \delta^{\prime}=0.006$ $\Rightarrow\ \gamma_{\!e}=0.0110,\ \delta_{\!e}=0.0165$. The blockwise invariants (for [[1a,a],[b,1b]][[1-a,a],[b,1-b]]) are
@@
D.4 Minimal Python excerpt (reproducibility).
@@
# GC block (asymmetric, to place y* $\approx$ 0.60)